# Leveraging partial stragglers within gradient coding

### v1

@@ I Introduction
The usage of the partial work performed by stragglers has been considered [10, 24, 25, 26, 27, 28] only within exact GC (i.e., approximate GC not considered); some of these apply in the communication-efficient setting. However, these approaches use multiple messages from the workers to the PS and thus incur a higher communication cost per iteration.
@@
Motivation: Consider Fig. 1(a) where the labels indicate the encoded gradients. Note that under ideal operating conditions when each worker operates at the same speed, the overall gradient can be computed as long as each $W_{i}$ processes its first chunk $\mathcal{D}_{i}$ for $i=1,\dots,3$. Thus, it is quite wasteful for the workers to continue processing their second assigned chunk as specified in the original GC scheme; it would double the computation time. In addition, the original GC formulation ignores partial work by slow but not failed workers; in the sequel, we refer to these as “partial stragglers”. For instance, in Fig. 1(a), we consider a scenario, where $W_{1}$ is slow and $W_{3}$ is failed. The state of the computation is such that there is enough information for the PS to obtain the full gradient. However, under the original GC model, $W_{1}$ will either wait until it processes $\mathcal{D}_{2}$ before generating the encoded gradient to be sent to the PS, or the PS will treat $W_{1}$ as failed and will have to settle for an approximate gradient.
Gradient coding can be viewed as an application of coding-theoretic techniques [29] to distributed learning; it allows the PS to be resilient to failures with very little coordination in the cluster. We note here that within classical coding theory [29] most construction of erasure codes do not consider feedback from the receiver to the sender since such feedback may be expensive or noisy when the sender and receiver are at remote locations or not even necessary. However, feedback is quite easy to implement in the distributed learning setup.
@@ II Gradient Coding for partial stragglers
Following this, the workers need to decide their own encoding coefficients for the gradients that they have computed. Each gradient $g_{\mathcal{D}_{i}}$ is block-partitioned into $\ell$ disjoint parts $g_{\mathcal{D}_{i}}[k],k=0,\dots,\ell-1$. Each worker forms a matrix $\mathbf{B}$ of dimension $m\times N\ell$ that consists of indeterminates that specify the encoding coefficients of all the workers (see example in upcoming Section II-B). Matrix $\mathbf{B}$ consists of $N$ block-columns, where each block-column itself consists of $\ell$ columns of length-$m$, i.e., $\mathbf{B}=[\mathbf{B}^{(1)}~{}|~{}\mathbf{B}^{(2)}~{}|~{}\dots~{}|~{}\mathbf{B}^{(N)}]$. The $j$-th block column, $\mathbf{B}^{(j)}$ is associated with $g_{\mathcal{D}_{j}}[k],k=0,\dots,\ell-1$. Based on the global state vector $\psi$, all workers know whether a chunk $\mathcal{D}_{j},j\in[N]$ has been processed by worker $W_{i},i\in[m]$. If $\mathcal{D}_{j}$ has been processed by $W_{i}$, then the $i$-th row of $\mathbf{B}^{(j)}$ is populated with indeterminates, otherwise these entries are set to zero. Once the indeterminates are found (see Algorithm 1 for a description), worker $W_{\beta}$ encodes its gradients as
@@
$W_{j}$ then solves a minimum-$\ell_{2}$-norm least-squares solution to determine its encoding coefficients (see (5)). This ensures the solution is unique [30] even if the corresponding problem is under-determined. Thus, the workers automatically agree on the same solution.
@@ II-A Analysis of Algorithm 1
Approximate GC analysis: It is possible that the PS sends the encode-and-transmit vector when some $\mathcal{D}_{i}$ has been processed $\Delta\leq\ell-1$ times. In this case, the corresponding step for $\mathbf{B}^{(i)}$ in (5) will be an over-determined least-squares procedure, which implies that there will be a non-zero error associated with it. Let $X$ be the relevant ×$\ell\times\Delta$ submatrix of $\mathbf{R}$ where we have $\Delta<\ell$ now, and recall that all entries of $X$ are i.i.d. $N(0,1)$ random variables. The squared error corresponding to $\mathbf{B}^{(i)}$ can be expressed as $\sum_{i=0}^{\ell-1}||XX^{\dagger}\vec{e}_{i}-\vec{e}_{i}||_{2}^{2}$ ($X^{\dagger}$ denotes the pseudo-inverse [30]). Let $X=USV^{T}$ denote the SVD of $X$, where $U$ and $V$ are orthogonal matrices of dimension $\ell\times\ell$ and $\Delta\times\Delta$ respectively and $S=[D~{}|~{}0]^{T}$ where $D$ is $\Delta\times\Delta$ matrix with positive entries on the diagonal, and 00 represents a Δ×($\Delta\times(\ell-\Delta)$ matrix of zeros. Then, $X^{\dagger}=V[D^{-1}~{}|~{}0]U^{T}$. It is well known ([32], Remark 5.2.8) that for a matrix with i.i.d. $N(0,1)$ entries, the singular vectors are uniformly distributed on the unit-sphere. Therefore, the expected squared error becomes
@@
Complexity analysis: The time complexity of each least-squares problem is $O(\Delta^{2}\ell)$ [33] and the $i$-th worker solves at most $\delta_{i}$ of them independently and in parallel. The marginal cost of this calculation as against the calculation of the actual gradients will be very small in most practical settings.
@@ II-B Illustrative Example
In this example, $W_{5}$ has processed $\mathcal{D}_{1},\mathcal{D}_{4},\mathcal{D}_{5}$ so that $\tilde{B}_{5}=[\mathbf{B}^{(1)}~{}\mathbf{B}^{(4)}~{}\mathbf{B}^{(5)}]$. Note that the PS requires the vectors $[\vec{e}_{0}^{T}~{}\vec{e}_{0}^{T}~{}\vec{e}_{0}^{T}~{}\vec{e}_{0}^{T}~{}\vec{e}_{0}^{T}]^{T}$ and $[\vec{e}_{1}^{T}~{}\vec{e}_{1}^{T}~{}\vec{e}_{1}^{T}~{}\vec{e}_{1}^{T}~{}\vec{e}_{1}^{T}]^{T}$ to lie in the row-space of $\mathbf{B}$ so that it can recover $\sum_{i=1}^{5}g_{\mathcal{D}_{i}}[k],k=0,1$ from the encoded gradients. Towards this end, e.g., $W_{5}$ solves (5) in Algorithm 1 with the matrix $\tilde{B}_{5}$, i.e.,
@@ III Chunk Ordering Algorithm
We work instead with a combinatorial metric that depends on the total number of chunks that the cluster has to process in the worst case, such that at least a single copy of each $\mathcal{D}_{i}$ is processed; this was also used in [34, 35] in a coded matrix computation context. In particular, for a given $\mathcal{D}_{i}$, let $Q_{i}$ denote the maximum number of chunks that can be processed by the cluster such that none of the copies of $\mathcal{D}_{i}$ are processed (see Figs. 2(a) and 2(b)) and let $Q_{\max}=\max_{i=1,\dots,N}Q_{i}$. Thus, $1+Q_{\max}$ is a metric that quantifies the worst-case number of chunks that need to be processed before at least a single copy of every $\mathcal{D}_{i}$ has been processed. Here, the worst-case is over the speeds of the different workers. We say that an assignment is optimal if it achieves the lowest possible $Q_{\max}$ for a given assignment matrix. Indeed, Figs. 2(a) and 2(b) show two different orderings for the same assignment matrix with different values of $Q_{\max}$.
@@ IV Numerical Experiments and Comparisons
Figure 3: (a) Mean-squared error (MSE)figu vs. $T$ for an approximate GC scenario. Blue curves: proposed protocol with $\ell=1,2,3$, purple curves: corresponding MSE estimates, and red curve: original GC protocol with $\ell=1$. Error bars correspond to the 95% confidence interval. (b) Completion time vs. $\ell$ for exact GC scenario with two different assignment matrices. Blue curves: proposed protocol, green curves: original GC protocol. Error bars correspond to one standard deviation.
Next, we present comparisons with the original GC protocol via a simulation of the distributed system for both the exact and approximate GC scenarios. These simulations were performed within Google Colab (2.20GHz processor, 13.2 GB RAM). In both scenarios, we simulated node failures and slow-downs. We generated a random vector of length-$m$ where $\alpha$ workers uniformly at random are chosen to be failed ($\alpha$ is chosen based on the scenario). For the other workers, the amount of time taken to process a chunk is chosen i.i.d. from an exponential distribution with parameter $\mu=1$. The entries of the matrix $\mathbf{R}$ are chosen i.i.d. from the standard normal $N(0,1)$.
Approximate GC simulation: We considered two different random regular graphs, $G_{i},i=1,2$ with sizes $m=100,150$ and degree $\nu=6$. These graphs have second eigenvalues: 4.29 and 4.3, i.e., less than $2\sqrt{\nu-1}$ so they can be considered as Ramanujan graphs [39] used in [6]. Their adjacency matrix was used as the assignment matrix. The number of failures $\alpha$ was set to $12\nu+$ 1.
@@
For the original GC protocol, we compute the least squares solution $\hat{r}$ for minimizing $||Ar-\mathds{1}||^{2}_{2}$. Here, $\hat{r}$ such that $\hat{r}_{i}=0$ if $W_{i}$ has not completed processing all its chunks at time $T$. For our proposed protocol, we leverage partial work completed by $T$, as described in Section II. The overall error is computed as the sum of the errors for the LS solution for each $z_{i}^{T},i=0,\dots,\ell-1$. Each data point on the curves was chosen by performing 100 simulations of failures and slow-downs. We also plot the expected value of the error, derived in (6).
@@
In Fig. 6 in Appendix B, we compare the performance of our approach under the optimal ordering (cf. Section III) and choosing a random ordering. Our optimal ordering, which can be found efficiently, clearly has a better performance.
Exact GC simulation: We considered (i) the cyclic assignment [5] with $N=m=100$ and $\delta=6$, and (ii) the assignment based on graph $G_{1}$ discussed above. The number of failures $\alpha$ in the simulations is set to $\delta-\ell$ so that exact gradient reconstruction is possible. For both approaches, we determined the time $T$ such that each chunk is processed at least $\ell$ times across the cluster (for original GC we only consider workers that have processed all their chunks). These values were averaged over 100 runs for each value of $\ell$. In Fig. 3(b) we clearly observe that the exact gradient can be computed using our method is approximately half the time as compared to the original GC protocol. We note here that when $\ell\geq 2$, the known techniques for exact GC based on Lagrange interpolation are numerically unstable. Thus, the gradient recovered using the original approach will in general not be useful.
Fig. 6 in Appendix B compares the completion times of random vs. optimal ordering; the optimal order is clearly better.
V Conclusions
@@ Appendix B Additional numerical experiments
Fig. 5 shows the mean-squared-error (MSE) comparing the original GC approach with $\ell=1$ and our partial GC approach for $\ell=1,2,3$ for graph $G_{2}$ with $m=150$ vertices. The results are similar in spirit to the results for case of $G_{1}$ that has 100 vertices. Namely, our MSE is orders of magnitude lower than the original GC approach, even when we consider $\ell\geq 2$.
Figure 5: Mean-squared error (MSE) vs. $T$ for an approximate GC scenario corresponding to an assignment matrix stemming from graph $G_{2}$. Error bars correspond to the 95% confidence interval. Blue curves: proposed protocol with $\ell=1,2,3$, purple curves: corresponding MSE estimates, and red curve: original GC protocol with $\ell=1$.
In Figs. 6(a) and 6(b) we study the impact of chunk ordering within the workers for the assignment matrices corresponding to graphs $G_{1}$ and $G_{2}$ respectively. Each data point on the curves corresponds to 100 simulations (setup described in Section IV). In particular, in Fig. 6(a) corresponding to the case of $\ell=3$ for $G_{1}$, we observe that the MSE of the optimal ordering consistently remains lower than the MSE of the random ordering and can in fact be multiple orders of magnitude lower when the encode-and-transmit signal is sent around $9\leq T\leq 10$. A similar pattern can be observed in Fig. 6(b), which shows the case of $\ell=3$ and $G_{2}$. We note here that we only selected a specific value of $\ell$ so that the plot can be visualized easily. Similar patterns and trends can be noticed even for other values of $\ell$ and other assignment matrices.
Figure 6: Mean-squared error (MSE) vs. $T$ for the approximate GC scenario when considering random chunk ordering and optimal chunk ordering within our protocol with $\ell=3$. Error bars correspond to the 95% confidence interval. (a) Assignment matrix corresponding to graph $G_{1}$. (a) Assignment matrix corresponding to graph $G_{2}$.
Fig. 7 shows the results of a similar experiment comparing the random chunk ordering and our optimal chunk ordering in terms of completion time for exact GC. The assignment matrix corresponds to the graph $G_{1}$ discussed in Section IV. Once again, each data point corresponds to 100 simulations (setup described in Section IV). The results indicate that in an average sense, the completion time of the optimal ordering is clearly lower.

### v2

@@ I Introduction
The usage of the partial work performed by stragglers has been considered [10, 24, 25, 26, 27, 28] only within exact GC (i.e., approximate GC has not been considered); some of these apply in the communication-efficient setting. However, these approaches use multiple messages from the workers to the PS and thus incur a higher communication cost per iteration.
@@
Motivation: Consider Fig. 1(a) where the edge labels indicate the encoded gradients. Note that under ideal operating conditions when each worker operates at the same speed, the overall gradient can be computed as long as each $W_{i}$ processes its first chunk $\mathcal{D}_{i}$ for $i=1,\dots,3$. Thus, it is quite wasteful for the workers to continue processing their second assigned chunk as specified in the original GC scheme; it would double the computation time. In addition, the original GC formulation ignores partial work by slow but not failed workers; in the sequel, we refer to these as “partial stragglers”. For instance, in Fig. 1(a), we consider a scenario, where $W_{1}$ is slow and $W_{3}$ is failed. The state of the computation is such that there is enough information for the PS to obtain the full gradient. However, under the original GC model, $W_{1}$ will either wait until it processes $\mathcal{D}_{2}$ before generating the encoded gradient to be sent to the PS, or the PS will treat $W_{1}$ as failed and will have to settle for an approximate gradient.
Gradient coding can be viewed as an application of coding-theoretic techniques [29] to distributed learning; it allows the PS to be resilient to failures with very little coordination in the cluster. We note here that within classical coding theory [29] most constructions of erasure codes do not consider feedback from the receiver to the sender since such feedback may be expensive or noisy when the sender and receiver are at remote locations or not even necessary ([30], Chap. 7). However, feedback is quite easy to implement in the distributed learning setup.
@@ II Gradient Coding for partial stragglers
Following this, the workers need to decide their own encoding coefficients for the gradients that they have computed. Each gradient $g_{\mathcal{D}_{i}}$ is block-partitioned into $\ell$ disjoint parts $g_{\mathcal{D}_{i}}[k],k=0,\dots,\ell-1$. Each worker forms a matrix $\mathbf{B}$ of dimension $m\times N\ell$ that consists of indeterminates that specify the encoding coefficients of all the workers (see example in the upcoming Section II-B). Matrix $\mathbf{B}$ consists of $N$ block-columns, where each block-column itself consists of $\ell$ columns of length-$m$, i.e., $\mathbf{B}=[\mathbf{B}^{(1)}~{}|~{}\mathbf{B}^{(2)}~{}|~{}\dots~{}|~{}\mathbf{B}^{(N)}]$. The $j$-th block column, $\mathbf{B}^{(j)}$ is associated with $g_{\mathcal{D}_{j}}[k],k=0,\dots,\ell-1$. Based on the global state vector $\psi$, all workers know whether a chunk $\mathcal{D}_{j},j\in[N]$ has been processed by worker $W_{i},i\in[m]$. If $\mathcal{D}_{j}$ has been processed by $W_{i}$, then the $i$-th row of $\mathbf{B}^{(j)}$ is populated with indeterminates, otherwise these entries are set to zero. Once the indeterminates are found (see Algorithm 1 for a description), worker $W_{\beta}$ encodes its gradients as
@@
$W_{j}$ then solves a minimum-$\ell_{2}$-norm least-squares solution to determine its encoding coefficients (see (5) in Algorithm 1). This ensures the solution is unique [31] even if the corresponding problem is under-determined. Thus, the workers automatically agree on the same solution.
@@
If $\ell=1$, then it is possible to arrive at a protocol whereby the workers agree to transmit appropriately weighted partial sums of their gradients, so that the PS can exactly/approximately recover the sum. However, in the communication-efficient setting when $\ell>1$, the encoding is no longer straightforward. Thus, $\ell>1$ is the main scenario we consider in what follows.
Remark 2.
@@ II-A Analysis of Algorithm 1
Approximate GC analysis: It is possible that the PS sends the encode-and-transmit vector when some $\mathcal{D}_{i}$ has been processed $\Delta\leq\ell-1$ times. In this case, the corresponding step for $\mathbf{B}^{(i)}$ in (5) will be an over-determined least-squares procedure, which implies that there will be a non-zero error associated with it. Let $X$ be the relevant ×$\ell\times\Delta$ submatrix of $\mathbf{R}$ where we have $\Delta<\ell$ now, and recall that all entries of $X$ are i.i.d. $N(0,1)$ random variables. The squared error corresponding to $\mathbf{B}^{(i)}$ can be expressed as $\sum_{i=0}^{\ell-1}||XX^{\dagger}\vec{e}_{i}-\vec{e}_{i}||_{2}^{2}$ ($X^{\dagger}$ denotes the pseudo-inverse [31]). Let $X=USV^{T}$ denote the SVD of $X$, where $U$ and $V$ are orthogonal matrices of dimension $\ell\times\ell$ and $\Delta\times\Delta$ respectively and $S=[D~{}|~{}0]^{T}$ where $D$ is a $\Delta\times\Delta$ matrix with positive entries on the diagonal, and 00 represents a Δ×($\Delta\times(\ell-\Delta)$ matrix of zeros. Then, $X^{\dagger}=V[D^{-1}~{}|~{}0]U^{T}$. It is well known ([33], Remark 5.2.8) that for a matrix with i.i.d. $N(0,1)$ entries, the singular vectors are uniformly distributed on the unit-sphere. Therefore, the expected squared error becomes
@@
Complexity analysis: The time complexity of each least-squares problem is $O(\Delta^{2}\ell)$ [34] and the $i$-th worker solves at most $\ell\delta_{i}$ of them independently and in parallel. The marginal cost of this calculation as against the calculation of the actual gradients will be very small in most practical settings.
@@ II-B Illustrative Example
In this example, $W_{5}$ has processed $\mathcal{D}_{1},\mathcal{D}_{4},\mathcal{D}_{5}$ so that $\tilde{B}_{5}=[\mathbf{B}^{(1)}~{}\mathbf{B}^{(4)}~{}\mathbf{B}^{(5)}]$. Note that the PS requires the vectors $[\vec{e}_{0}^{T}~{}\vec{e}_{0}^{T}~{}\vec{e}_{0}^{T}~{}\vec{e}_{0}^{T}~{}\vec{e}_{0}^{T}]$ and $[\vec{e}_{1}^{T}~{}\vec{e}_{1}^{T}~{}\vec{e}_{1}^{T}~{}\vec{e}_{1}^{T}~{}\vec{e}_{1}^{T}]$ to lie in the row-space of $\mathbf{B}$ so that it can recover $\sum_{i=1}^{5}g_{\mathcal{D}_{i}}[k],k=0,1$ from the encoded gradients. Towards this end, e.g., $W_{5}$ solves (5) in Algorithm 1 with the matrix $\tilde{B}_{5}$, i.e.,
@@ III Chunk Ordering Algorithm
We work instead with a combinatorial metric that depends on the total number of chunks that the cluster has to process in the worst case, such that at least a single copy of each $\mathcal{D}_{i}$ is processed; this was also used in [35, 36] in a coded matrix computation context. This metric minimizes the worst case number of chunks that the cluster needs to process in the case when $\ell=1$. In particular, for a given $\mathcal{D}_{i}$, let $Q_{i}$ denote the maximum number of chunks that can be processed by the cluster such that none of the copies of $\mathcal{D}_{i}$ are processed (see Figs. 2(a) and 2(b)) and let $Q_{\max}=\max_{i=1,\dots,N}Q_{i}$. Thus, $1+Q_{\max}$ is a metric that quantifies the worst-case number of chunks that need to be processed before at least a single copy of every $\mathcal{D}_{i}$ is guaranteed to be processed. Here, the worst-case is over the speeds of the different workers. We say that an assignment is optimal if it achieves the lowest possible $Q_{\max}$ for a given assignment matrix. Indeed, Figs. 2(a) and 2(b) show two different orderings for the same assignment matrix with different values of $Q_{\max}$.
@@
Remark 4.
The proposed algorithm above finds the optimal ordering when considering the case of $\ell=1$ with $N=m$ (number of chunks equal to number of workers); it only applies when $N=m$. While this algorithm is expected to have better performance than a randomly chosen ordering in the case of higher $\ell$, we do not have an optimal construction in this case.
@@ IV Numerical Experiments and Comparisons
Figure 3: (a) Mean-squared error (MSE) vs. $T$ for an approximate GC scenario. Blue curves: proposed protocol with $\ell=1,2,3$, purple curves: corresponding MSE estimates, and red curve: original GC protocol with $\ell=1$. Error bars correspond to one standard deviation. (b) Completion time vs. $\ell$ for exact GC scenario with two different assignment matrices. Blue curves: proposed protocol, green curves: original GC protocol. Error bars correspond to one standard deviation.
In what follows, we present comparisons with the original GC protocol via a simulation of the distributed system for both the exact and approximate GC scenarios. All software code for recreating these results can be found at [40]. These simulations were performed within a MacBook Pro (M1 chip, 16 GB RAM). In both scenarios, we simulated node failures and slow-downs. We generated a random vector of length-$m$ where $\alpha$ workers uniformly at random are chosen to be failed ($\alpha$ is chosen based on the scenario). For the other workers, the amount of time taken to process a chunk is chosen i.i.d. from an exponential distribution with parameter $\mu=1$. The entries of the matrix $\mathbf{R}$ are chosen i.i.d. from the standard normal $N(0,1)$.
Approximate GC simulation: We considered two different random regular graphs, $G_{i},i=1,2$ with sizes $m=200,300$ and degree $\nu=8$. These graphs have second-largest absolute value of eigenvalues: 5.14 and 5.23, i.e., less than $2\sqrt{\nu-1}$ so they can be considered as Ramanujan graphs [41] (these were used in [6]). Their adjacency matrices were used as the assignment matrix. The number of failures $\alpha$ was set to $\nu-1$.
@@
For the original GC protocol, we compute the least squares solution $\hat{r}$ for minimizing $||Ar-\mathds{1}||^{2}_{2}$. Here, $\hat{r}$ such that $\hat{r}_{i}=0$ if $W_{i}$ has not completed processing all its chunks at time $T$. For our proposed protocol, we leverage partial work completed by $T$, as described in Section II. The overall error is computed as the sum of the errors for the least-squares solution for each $z_{i}^{T},i=0,\dots,\ell-1$. Each data point on the curves was chosen by performing 1000 simulations of failures and slow-downs. We have also plotted the expected value of the error, derived in (6).
@@
In Fig. 6 in Appendix B, we compare the performance of our approach under the optimal ordering (cf. Section III) and an appropriately chosen random ordering. The random ordering is picked as follows. We generated 100 independent random orderings and selected the one with the best (smallest) $Q_{\max}(O)$. Our optimal ordering, which can be found efficiently, clearly has a better performance.
Exact GC simulation: We considered (i) the cyclic assignment [5] with $N=m=200$ and $\delta=8$, and (ii) the assignment based on graph $G_{1}$ discussed above. The number of failures $\alpha$ in the simulations is set to $\delta-\ell$ so that exact gradient reconstruction is possible. For both approaches, we determined the time $T$ such that each chunk is processed at least $\ell$ times across the cluster (for original GC we only consider workers that have processed all their chunks). These values were averaged over 1000 runs for each value of $\ell$. In Fig. 3(b) we clearly observe that the exact gradient can be computed using our method is approximately half the time as compared to the original GC protocol. We note that the average completion time for the original GC protocol ($G_{1}$-based assignment) for $\ell=1$ is about $T=6$ time units. However, the MSE for the original GC protocol in the approximate scenario continues to be high at $T=24$. The reason is that there are $\nu-1$ failures that are introduced in the simulation. The approximate GC recovery operates by solving a least-squares problem for the fixed $G_{1}$-based assignment. Thus, the MSE does not necessarily drop to zero even if one copy of each chunk has been processed in the cluster. However, there are exact recovery algorithms that one can use in this case. We note here that when $\ell\geq 2$, the known techniques for exact recovery for the original GC protocol are based on Lagrange interpolation and are numerically unstable. Thus, the gradient recovered using the original approach will in general not be useful.
Fig. 7 in Appendix B compares the completion times of random vs. optimal ordering; the optimal ordering is clearly better.
V Limitations of our work
Our chunk ordering algorithm (Section III) is optimal only in the case when the assignment matrix has $N=m$ and $\gamma_{i}=\delta_{i}=\delta$ for $i\in[m]$. While several well known gradient coding schemes, especially those from regular graphs, satisfy this property, there are others that do not. Our results in Section IV with $\ell\geq 2$ are simulations of the actual distributed cluster and indicate lower MSE than the original GC protocol. However, we do not have actual cloud platform statistics on the reduction in communication time within our method. We do however expect this reduction to be quite significant.
VI Conclusions
@@ Appendix B Additional numerical experiments
Fig. 5 shows the mean-squared-error (MSE) comparing the original GC approach with $\ell=1$ and our partial GC approach for $\ell=1,2,3$ for graph $G_{2}$ with $m=300$ vertices. The results are similar in spirit to the results for the case of $G_{1}$ that has 200 vertices. Namely, our MSE is orders of magnitude lower than the original GC approach, even when we consider $\ell\geq 2$.
Figure 5: Mean-squared error (MSE) vs. $T$ for an approximate GC scenario corresponding to an assignment matrix stemming from graph $G_{2}$. Error bars correspond to one standard deviation. Blue curves: proposed protocol with $\ell=1,2,3$, purple curves: corresponding MSE estimates, and red curve: original GC protocol with $\ell=1$.
In Figs. 6(a) and 6(b) we study the impact of chunk ordering within the workers for the assignment matrices corresponding to graphs $G_{1}$ and $G_{2}$ respectively. Each data point on the curves corresponds to 1000 simulations (setup described in Section IV). In particular, in Fig. 6(a) corresponding to the case of $\ell=1,2,3$ for $G_{1}$, we observe that the MSE of the optimal ordering consistently remains lower than the MSE of the random ordering and can in fact be multiple orders of magnitude lower when the encode-and-transmit signal is sent at certain time intervals. A similar pattern can be observed in Fig. 6(b), which shows the case of $\ell=1,2,3$ and $G_{2}$.
Figure 6: Mean-squared error (MSE) vs. $T$ for the approximate GC scenario when considering random chunk ordering and optimal chunk ordering within our protocol with $\ell=1,2,3$. Error bars correspond to one standard deviation. (a) Assignment matrix corresponding to graph $G_{1}$. (a) Assignment matrix corresponding to graph $G_{2}$.
Fig. 7 shows the results of a similar experiment comparing the random chunk ordering and our optimal chunk ordering in terms of completion time for exact GC. The assignment matrix corresponds to the graph $G_{1}$ discussed in Section IV. The random ordering is chosen by sampling 100 independent random orderings and selecting one with the smallest $Q_{max}(O)$. A data point is then generated by running 1000 simulations with the selected random ordering. The results indicate that in an average sense, the completion time of the optimal ordering is clearly lower.